# Gaussian Splashing: Dynamic Fluid Synthesis with Gaussian Splatting

### v1

@@ 4.3. Position Based Fluids
We employ the position-based fluids (PBF) (Macklin and Müller, 2013) as our Lagrangian fluid synthesizer. To enforce the fluid incompressibility, PBF imposes a density constraint $C^{\rho}_{i}$ on each particle, maintaining the integrated density $\rho_{i}$ computed by the SPH kernel as:
(9) $C_{i}^{\rho}=\frac{\rho_{i}}{\rho_{0}}-1=\sum_{j}\frac{m_{j}}{\rho_{0}}W(\bm{p}_{i}-\bm{p}_{j})-1,$
where $m_{j}$ is the mass of particle $j$. $\bm{p}_{i}$ is the position of particle $i$, and $W$ is the SPH kernel function. The constraint Jacobian is:
(10) $\nabla_{\bm{p}_{k}}C_{i}^{\rho}=\begin{cases}\displaystyle\sum_{j}\frac{m_{j}}{\rho_{0}}\nabla_{\bm{p}_{i}}W(\bm{p}_{i}-\bm{p}_{j}),&k=i\\ \displaystyle\frac{m_{j}}{\rho_{0}}\nabla_{\bm{p}_{j}}W(\bm{p}_{i}-\bm{p}_{j}),&k=j.\end{cases}$
Figure 4. Detection of surface particles.  An interior particle is detected if its screen is widely shadowed by its neighbors. A boundary particle is detected if at least one part of the particle’s screen is not shadowed.
@@
GSP also includes a position-based tension model (Xing et al., 2022) to better capture the dynamics of the fluid surface. We first detect whether a particle (i.e., a Gaussian kernel) is on the fluid surface based on occlusion estimation. Specifically, we encapsulate a particle with a spherical cover or screen. Each of its neighboring particles generates a projection on the screen (because a particle has a finite volume). The particle is considered on the fluid surface if the total projection area from its neighbors is below a given threshold. Tensions tends to minimize surface area. Therefore, PBF applies an area constraint to each surface particle to minimize the local surface area nearby. We start by calculating the normal $\bm{n}_{i}$ of surface particles $i$ as:
(11) $\bm{n}_{i}=\text{normalize}(-\nabla_{\bm{p}_{i}}C_{i}^{\rho}),$
@@
(12) $C_{i}^{A}=\sum_{t\in T(i)}\frac{1}{2}\left\|(\bm{p}_{t^{2}}-\bm{p}_{t^{1}})\times(\bm{p}_{t^{3}}-\bm{p}_{t^{1}})\right\|$
(13) $C_{ij}^{D}={\min}\left\{0,\left\|\bm{p}_{i}-\bm{p}_{j}\right\|-d_{0}\right\},$
where $d_{0}$ is the distance threshold. The Jacobian of these constraints are:
(14) $\nabla_{t^{1}}C_{t}^{A}(\bm{p})$ $=\frac{1}{2}\frac{(\bm{p}_{t^{2}}-\bm{p}_{t^{1}})\times(\bm{p}_{t^{3}}-\bm{p}_{t^{1}})}{\left\|(\bm{p}_{t^{2}}-\bm{p}_{t^{1}})\times(\bm{p}_{t^{3}}-\bm{p}_{t^{1}})\right\|}\times(\bm{p}_{t^{3}}-\bm{p}_{t^{2}}),$
$\nabla_{t^{2}}C_{t}^{A}(\bm{p})$ $=\frac{1}{2}\frac{(\bm{p}_{t^{3}}-\bm{p}_{t^{2}})\times(\bm{p}_{t^{1}}-\bm{p}_{t^{2}})}{\left\|(\bm{p}_{t^{3}}-\bm{p}_{t^{2}})\times(\bm{p}_{t^{1}}-\bm{p}_{t^{2}})\right\|}\times(\bm{p}_{t^{1}}-\bm{p}_{t^{3}}),$
$\nabla_{t^{3}}C_{t}^{A}(\bm{p})$ $=\frac{1}{2}\frac{(\bm{p}_{t^{1}}-\bm{p}_{t^{3}})\times(\bm{p}_{t^{2}}-\bm{p}_{t^{3}})}{\left\|(\bm{p}_{t^{1}}-\bm{p}_{t^{3}})\times(\bm{p}_{t^{2}}-\bm{p}_{t^{3}})\right\|}\times(\bm{p}_{t^{2}}-\bm{p}_{t^{1}}),$
@@
4.4. Rendering
(15) $\bm{A}_{p}^{t}=\bm{F}_{p}\bm{A}_{p}\bm{F}_{p}^{\top},\;\text{and}\;\bm{n}_{p}^{t}=\frac{\bm{F}_{p}^{-\top}\bm{n}_{p}}{\left\|\bm{F}_{p}^{-\top}\bm{n}_{p}\right\|}.$
We then shade the deformed Gaussian kernels $\{\bm{x}^{t}_{p},\sigma_{p},\bm{A}_{p}^{t},\bm{n}_{p}^{t}\}$ with material $\{\bm{d}_{p},s_{p},\rho_{p}\}$ i.e., Eq. (5) and splat them into an image $\bm{c}^{\rm bg}$. The dynamic particle-based fluids are rendered with ellipsoids splatting (Macklin and Müller, 2013), which is inherently compatible with the existing 3DGS pipeline. We begin by generating fluid Gaussian kernels at each fluid particle. The initial covariance $\bm{A}_{p}$ of each kernel is determined by the particle radius. The normal $\bm{n}_{p}$ adopts the surface normal of the nearest surface fluid particle from PBF simulation. We then align the minimal axis of the Gaussian kernel with $\bm{n}_{p}$ by a single rotation and correspondingly adjust the initial covariance.
We proceed to set up the appearance material for each fluid Gaussian kernel. It is known that fluid’s appearance exhibits strong reflection and refraction effects. We employ the current PBR workflow to model the reflection, which adopts the same formula of Eq. (5). We set a specular material ($s_{p}=1$, $\rho_{p}=0.05$ in our experiments) for all fluid kernels to imitate reflective behavior. The refraction needs careful treatment, which we model into a thickness-dependent diffuse color $\bm{d}_{p}$. As the fluid thickness increases, the absorption of light within the fluid intensifies, resulting in reduced visibility for objects behind. Conversely, when there is less fluid present, it exhibits a more transparent appearance. The fluid thickness $\tau$ comes from the modified splatting pipeline, with the alpha blending replaced by additive blending. The refraction color $\bm{d}_{p}$ is then represented by Beer’s Law (Swinehart, 1962):
(16) $\bm{d}_{p}=e^{-k\tau_{p}}\bm{c}_{p}^{\rm bg}.$
Here, the absorption coefficient $k$ is defined differently for each color channel, $\tau_{p}$,$\bm{c}_{p}^{\rm bg}$ is the fluid thickness and background back-projected to each Gaussian kernel respectively. Note that for background back-projection, a distortion $\beta\bm{n}_{p}$ is added to mimic the change of light path. Opacity $\sigma_{p}$ is set to $1$ as most of transmission and refraction has already been modeled into $\bm{d}_{p}$. We finally shade all fluid Gaussian kernels $\{\bm{x}^{t},\sigma_{p},\bm{A}_{p}^{t},\bm{n}_{p}^{t}\,\bm{d}_{p},s_{p},\rho_{p}\}$ and splat them to the fluid rendering result $\bm{c}^{\rm fluid}$ with Gaussian splatting. The final rendering result is achieved by combining the $\bm{c}^{\rm bg}$ and $\bm{c}^{\rm fluid}$, as shown in Fig. 5.
Figure 5. GSP rendering.  GSP synthesizes high-quality images corresponding to dynamically interacting fluids and solids. (a) The final rendered image combining rendered solids and fluids; (b) The rendering result of deforming solids; (c) The fluid thickness by additive splatting, where the darker color indicates the higher thickness; (d) The rendered dynamic fluids which is not occluded by solids.
4.5. Inpainting
Displaced object exposes unseen areas that were originally covered to the camera. Since they are not present in the input image, 3DGS is unable to recover the color and texture information of these areas, leading to black smudges and dirty textures in the result. GSP remedies this issue using an inpainting trick. First, we remove all the Gaussian kernels of the object that may be displaced. We then use LaMa (Suvorov et al., 2022) to inpaint the rendered results, noting the new colors of pixels originally in spots and assigning them to the first Gaussian kernel encountered by the rays emitted from these pixels. We average the recorded colors on all noted Gaussian kernels for their diffuse color, set their opacity to $1$, and their specular color to $\bm{0}$ to prevent highlights. The inpainted 3DGS rendering becomes much improved, as shown in Fig. 6.
Figure 6. 3DGS inpainting.  In this indoor scene, both the paper cup and the stuffed toy dog are segmented from the input image (left). 3DGS leaves empty spots and dirty textures blended from irrelevant kernels, as highlighted in the middle figure. Applying the inpainting with generative AI (Suvorov et al., 2022) ameliorates this issue (right).
5. Experiments
Table 1. Time performance.  We report detailed time statistics for experiments reported in the paper. NeuS gives the total time of mesh reconstruction. Train is the total time of training 3DGS. Both of the above measurements are in hours. However, all subsequent time-related evaluations are presented in seconds. #normal-#\## BG, #normal-#\## Solids and #normal-#\## Fluids represent the number of Gaussian kernels for the background, the foreground solid, and the fluid, respectively. Sim is the total time of simulation per time step. Tension represents the time taken for each surface particle detection. Render represents the total rendering time. Render BG, Render solids and Render fluids indicate the time spent on rendering the background, solids and fluids. Some experiments incorporates some specific settings, In the Book & Lego case, we only account for the simulation and rendering performance after the book transforms into water and the Lego excavator falls. Train is the sum of the book and the excavator. #normal-#\## Fluids is the counts of Gaussian kernels of the book, and #normal-#\## solids is the counts of Gaussian kernels of the excavator. In the Can, since the solid is completely fixed, only the fluid is involved in the Sim.
Scene NeuS Train ##\## BG ##\## Solids ##\## Fluids Sim Tension Render BG Render solids Render fluids Render
Chair
(Fig. 9)
[…]
Garden
(Fig. 10)
[…]
Book & Lego
(Fig. 11)
[…]
Can
(Fig. 12)
[…]
Cup & dog
(Fig. 13)
[…]
Astronaut
(Fig. 14)
[…]
Figure 7. Anistropy regularization.  Anistropy regularization effectively maintains rendering quality under large deformations. Without the regularization term, 3DGS tends to generate fuzzy and spiky artifacts, especially near the surface of the model (left). When the regularization is applied, image quality is greatly improved with correct specular effects.
We implemented GSP pipeline using Python, C++ and CUDA on a desktop PC equipped with a 12-core Intel i7-12700F CPU and an NVIDIA RTX 3090 GPU. Specifically, for rendering part, we ported the published implementation of Gaussian shader (Jiang et al., 2023) and integrated our fluid rendering using PyTorch (Imambi et al., 2021). PBD/PBF engine was implemented with CUDA, and we also group independent constraints int colors to efficiently parallelize the constraint projections on the GPU. We set the simulation time step as $0.005$ seconds throughout the simulations. In our PBD solver, we used $10$ iterations for fluids and $50$ iterations for solids for our experiments, since mass particles on the solid models are stronger coupled than the ones in the fluid. During the PBF simulation, the surface particles of fluids are updated every two time steps. 3DGS is originally designed for view synthesis. 3DGS obtained from a static scene produces low-quality renders when Gaussian kernels undergo large rotational deformations. The anisotropy regularization is effective against this limitation as shown Fig. 7. Detailed statistics regarding the experiment settings and timings are reported in Tab. 1. Most experiments can also be found in the supplementary video.
Refer to caption Refer to caption
Figure 8. Ablation study of specular.  We demonstrate the impact of specular highlights on the quality of rendering. On the left is a fluid rendered with diffuse color only; On the right, surface reflective specular are added, which exhibits a more realistic and dynamic fluid.
To show the importance of specular highlights in fluid rendering, we show a side-by-side comparison in a 3DGS scene. As shown in Fig. 8, we compare fluids rendered using purely diffuse materials with those that incorporated specular reflections. The fluids without specular reflection appear almost smoke-like, while the inclusion of specular term significantly enhances the realism of the fluids.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@ 5. Experiments
Figure 9. Chair.  A soft chair fell into the pool, causing deformation and ripples.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Figure 10. Flooding garden.  Waters leak into the garden and submerge the table. As the water level goes up, the surface gets more vibrant and washes the potted plant away.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 11. From books to pool.  Three books stack on the desk. They are isolated and segmented in GSP. The user switches their solid Gaussians to fluid Gaussians to make an indoor pool and drops a LEGO excavator into it.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 12. Water droplets on can.  Droplets of water fall onto the surface of a soda can, coalesce due to surface tension and gradually overflow.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Figure 13. “Everything is water”.  Pouring water into the paper cup on the table and transforming the cup and a dog toy into water. The water spills out.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Figure 14. Black magic.  An astronaut in the space strucked by the black magic of the Trisolarans, and get transformed into a water sphere.
We test GSP in a diverse set of experiments. A “warm-up” test is shown in Fig. 9, where a soft chair from NeRF synthetic datasets (Mildenhall et al., 2020) falls into a pool, illustrating the two-way coupling between deformable solids and fluids. The chair undergoes deformation, floats due to buoyancy, and creates fluid ripples as a result. Fig. 10 showcases another fluid-solid interaction test. The garden scene is sourced from the Mip-NeRF 360 (Barron et al., 2022) dataset. The foreground objects consist of a fixed table and a potted plant. We pour water into the garden slowly. The water rises up and eventually sinks the table, and sweeps the plant.
GSP has a semantic segmentation module. Therefore, the user is able to freely manipulate the models in the scene. Furthermore, since everything is represented as Gaussian particles, GSP allows the user to transform the state of the model. An example is shown in Fig. 11. In this indoor scene, three stacking books are isolated and segmented before training the 3DGS. The user transforms them from solid objects to fluid and instantly creates an in-door pool. A pre-trained Lego excavator falls into the “pool”. It then floats due to the buoyancy. Another similar experiment is reported in Fig. 13. The scene includes a round white table, on which a paper cup and a stuffed toy dog are placed. Water is poured into the cup. The state of toy dog and the cup are changed to water, and splashes on the desk. As mentioned, we use LaMA to inpaint the table texture so that the user does not observe rendering artifacts when the liquefied cup and toy dog splashes away.
Our fluid simulator can also capture the surface tension within the PBD framework. This enables realistic low-volume fluid-solid interaction. As shown in Fig. 12, droplets of water continuously drip onto the top of the can until they reaches the capacity and overflow. The surface tension of the liquid causes the droplets to gradually aggregate at the surface. As more droplets fall onto the liquid surface, they rise above the edge of the can. Eventually, the accumulated water exceeds the limit of surface tension and spills over, resulting in overflow.
Fig. 14 showcases another interesting use of GSP, where the user applies Trisolarans’ black magic on an astronaut. The astronaut is strucked by the magic and is transformed to water. It finally collapses into a water ball in a zero-gravity space due to the presence of surface tension.
6. Conclusion
Gaussian Splashing is a novel pipeline combining versatile position-based dynamics with 3DGS. The principle design philosophy of GSP is to harness the consistency of volume particle-based discretization to enable integrated processing of various 3D graphics and vision tasks, such as 3D reconstruction, deformable simulation, fluid dynamics, and rendering from new camera angles. While the concept is straightforward, building GSP involves significant research and engineering efforts. The presence of fluid complicates the 3DGS processing due to the specular highlights at the fluid surface; Fluid-solid coupling resorts to accurate surface information; Large deformation on the solid object generates defective rendering; Displaced models also leave empty regions that the input images fail to capture. We overcome those difficulties by systematically integrating and adapting a collection of state-of-the-art technologies into the framework. As a result, GSP enables realistic view synthesis not only under novel camera poses but also with novel physically-based fluid/solid dynamics or even novel object state transform. It should be noted that incorporating physically-based fluid dynamics in NeRF/3DGS has not been explored previously. The primary contribution of this work is to showcase the feasibility of building a unified framework for integrated physics and learning-based 3D reconstruction. GSP still has many limitations. For instance, PBD is known to be less physically accurate. It may be worth generalizing PBD with other meshless simulation methods. The fluid rendering in GSP in its current form is far from perfect – ellipsoid splatting is an ideal candidate for position-based fluid but does not physically handle refraction. PBF may need a large number of fluid particles, which negatively impacts the rendering efficiency.

### v2

@@ 7.2 Position-Based Fluids
We employ the Position-Based Fluids (PBF) [40] as our Lagrangian fluid synthesizer. PBF is based on PBD, which means it also use constraint projections to simulate fluid behaviour. In PBF, fluid is composed of a large amount of particles. To enforce the fluid incompressibility, PBF imposes a density constraint $C^{\rho}_{i}$ on each particle, maintaining the integrated density $\rho_{i}$ computed by the SPH kernel as:
$C_{i}^{\rho}=\frac{\rho_{i}}{\rho_{0}}-1=\sum_{j}\frac{m_{j}}{\rho_{0}}W(\bm{p}_{i}-\bm{p}_{j},r)-1,$ (16)
where $m_{j}$ is the mass of particle $j$. $\bm{p}_{i}$ is the position of particle $i$, $W$ is the SPH kernel function and $r$ is the kernel radius. Intuitively, projecting this constraint to 00 ensures that the density at the current time remains consistent with the initial state. We use the following cubic SPH kernel:
$W(\bm{p},r)=\begin{cases}\frac{8}{\pi r^{3}}(6q^{2}(q-1)+1),&0\leq q\leq 0.5\\ \frac{16}{\pi r^{3}}(1-q)^{3},&0.5<q\leq 1\\ 0,&{\rm otherwise}\end{cases}$ (17)
where $q=\frac{\|\bm{p}\|}{r}$. The Jacobian of constraint is computed as:
$\nabla_{\bm{p}_{k}}C_{i}^{\rho}=\begin{cases}\displaystyle\sum_{j}\frac{m_{j}}{\rho_{0}}\nabla_{\bm{p}_{i}}W(\bm{p}_{i}-\bm{p}_{j},r),&k=i\\ \displaystyle\frac{m_{j}}{\rho_{0}}\nabla_{\bm{p}_{j}}W(\bm{p}_{i}-\bm{p}_{j},r),&k=j.\end{cases}$ (18)
The gradient of the kernel function is:
$\nabla_{\bm{p}}W(\bm{p},r)=\begin{cases}\frac{48}{\pi r^{5}}(3q-2)\bm{p},&0\leq\frac{\|\bm{p}\|}{r}\leq 0.5\\ -\frac{48}{\pi r^{5}}\frac{(1-q)^{2}}{q}\bm{p},&0.5<\frac{\|\bm{p}\|}{r}\leq 1\\ 0,&{\rm otherwise}\end{cases}$ (19)
@@
GSP also includes a position-based surface tension model [70] to better capture the dynamics of the fluid surface. We first detect whether a particle (i.e., a Gaussian kernel) is on the fluid surface based on occlusion estimation. Specifically, we encapsulate a particle with a spherical cover or screen. Each of its neighboring particles generates a projection on the screen (because a particle has a finite volume). The particle is considered on the fluid surface if the total projection area from its neighbors is below a given threshold.
In the original paper [70], surface detection is implemented on the CPU. It is noteworthy that surface detection can be parallelized on the GPU to expedite the simulation, as the calculation of each particle’s occluding ratio on the screen is independent of the others.
Figure 10: Detection of surface particles.  An interior particle is detected if its screen is widely shadowed by its neighbors. A boundary particle is detected if at least one part of the particle’s screen is not shadowed.
For each neighboring particle, its occluding area on the spherical screen is calculated as follows:
$\theta$ $={\rm tan^{-1}}(\frac{\Delta\bm{p}_{y}}{\Delta\bm{p}_{x}+\Delta\bm{p}_{z}^{2}})$
$\phi$ $={\rm tan^{-1}}(\frac{\Delta\bm{p}_{x}}{\Delta\bm{p}_{z}})$ (20)
$\Delta\theta$ $={\rm tan^{-1}}\frac{R}{\|\Delta p\|^{2}-R^{2}}$
$\Delta\phi$ $=\Delta\theta$
where $\Delta\bm{p}$ is the vector from the detection particle to the neighboring particle and $R$ is the particle radius. The shadowed area on the spherical screen is then $[\theta-\Delta\theta,\theta+\Delta\theta]\times[\phi-\Delta\phi,\phi+\Delta\phi]$. We parameterize the screen as an $18\times 36$ environment map, with each column of the environment map corresponding to 18 bits of an integer. We then mask 36 integers and count the mask ratio.
After detecting surface particles on the fluid, we apply tension on the surface. Tensions tends to minimize surface area. Therefore, PBF applies an area constraint to each surface particle to minimize the local surface area nearby. We start by calculating the normal $\bm{n}_{i}$ of surface particles $i$ as:
$\bm{n}_{i}=\text{normalize}(-\nabla_{\bm{p}_{i}}C_{i}^{\rho}),$ (21)
@@
$C_{i}^{A}=\sum_{t\in T(i)}\frac{1}{2}\left\|(\bm{p}_{t^{2}}-\bm{p}_{t^{1}})\times(\bm{p}_{t^{3}}-\bm{p}_{t^{1}})\right\|$ (22)
where $T(i)$ is the set of neighboring triangles for particle $i$. We use the 2D Delaunay triangulation to construct the triangles on the local surface. This process is sequential and cannot be parallelized on the GPU. However, it is sufficiently fast and we translate it from CPU to GPU.
Figure 11: Different sampling strategies.  We compare the results of different sampling strategies: (left) fill the particle based on the density grid calculated using Gaussian kernels [69], and (right) uniformly sample within NeuS reconstruction. The point distribution generated by vanilla 3DGS is uneven, which hardly samples the legs or seat of the chair.
To promote a more uniform particle distribution, additional distance constraints are introduced to push apart particles that are too close to each other:
$C_{ij}^{D}={\min}\left\{0,\left\|\bm{p}_{i}-\bm{p}_{j}\right\|-d_{0}\right\},$ (23)
where $d_{0}$ is the distance threshold. The Jacobian of aboved constraints are:
$\nabla_{t^{1}}C_{t}^{A}(\bm{p})$ $=\frac{(\bm{p}_{t^{2}}-\bm{p}_{t^{1}})\times(\bm{p}_{t^{3}}-\bm{p}_{t^{1}})\times(\bm{p}_{t^{3}}-\bm{p}_{t^{2}})}{2\left\|(\bm{p}_{t^{2}}-\bm{p}_{t^{1}})\times(\bm{p}_{t^{3}}-\bm{p}_{t^{1}})\right\|},$ (24)
$\nabla_{t^{2}}C_{t}^{A}(\bm{p})$ $=\frac{(\bm{p}_{t^{3}}-\bm{p}_{t^{2}})\times(\bm{p}_{t^{1}}-\bm{p}_{t^{2}})\times(\bm{p}_{t^{1}}-\bm{p}_{t^{3}})}{2\left\|(\bm{p}_{t^{3}}-\bm{p}_{t^{2}})\times(\bm{p}_{t^{1}}-\bm{p}_{t^{2}})\right\|},$
$\nabla_{t^{3}}C_{t}^{A}(\bm{p})$ $=\frac{(\bm{p}_{t^{1}}-\bm{p}_{t^{3}})\times(\bm{p}_{t^{2}}-\bm{p}_{t^{3}})\times(\bm{p}_{t^{2}}-\bm{p}_{t^{1}})}{2\left\|(\bm{p}_{t^{1}}-\bm{p}_{t^{3}})\times(\bm{p}_{t^{2}}-\bm{p}_{t^{3}})\right\|},$
@@
8 Sampling and Interpolation Details
In practice, we found that an uneven distribution of Gaussian kernels results in unstable and inaccurate motion synthesis, while a distribution that is too uniform can detrimentally affect rendering quality. Gaussian kernels tend to distribute primarily unevenly around the surfaces and edges of objects, leading to inaccurate boundary descriptions, which are crucial for interactions between objects in simulation. Conversely, adaptively distributed anisotropic Gaussian kernels are key to representing the spatially varying texture on the object. To address this issue, we maintain two separate sets of points: one sampled from the NeuS mesh surface for simulation, and the other consisting of trained Gaussian kernels for rendering. We compare the results of directly sampling from trained Gaussian kernels to those of sampling from NeuS in Fig. 11. The former method can result in sparsely sampled regions, especially for objects with thin parts, potentially affecting simulation quality.
We then discuss how to animate trained Gaussian kernels for rendering dynamics. Denote the set of trained Gaussian kernels for rendering as $S_{r}$, and the set of sampled points from NeuS for simulation as $S_{s}$. At time 00, we initialize GMLS kernels on $S_{s}$ and find the $k$ nearest neighbors $\{\bm{p}_{s,j}^{0}:j\in\mathcal{N}(i)\}$ from $S_{s}$ for each point $\bm{p_{r,i}}$ in $S_{r}$. Here, $\mathcal{N}(i)$ denotes the set of $k$ nearest neighboring particles’ indices of $\bm{p}_{r,i}$ in $S_{s}$ at time 00. As the simulation or motion synthesis proceeds, the position $\bm{p}^{n}_{s,j}$ evolves with time step $n$. We then update $\bm{p}_{r,j}$ by interpolating from $\{\bm{p}_{s,j}^{n}:j\in\mathcal{N}(i)\}$ with the pre-built GMLS kernel. The interpolation of deformation is achieved in the same way, replacing the physical quantity position $\bm{p}$ to deformation gradient $F$.
9 Rendering Details
Shadow
As shadows are crucial to the visual outcomes in dynamic scenes, we re-engineer nearly-soft shadows [14] into our system to enhance realism. Following shadow mapping, we splat all Gaussian kernels to a camera positioned at the point light’s location, which we denote as light-view. The point light is aligned with the direction of the significantly bright light in the environment map. The resolution of the light-view image is three times that of the original image resolution to address visual discrepancies caused by under-sampling.
We then reproject the points seen in the camera view to the light-view and compare their depths to the previously splatted light-view depth image. A larger depth indicates that the point is occluded by a nearer point and will therefore cast a shadow. A more robust variance method is discussed in [14]. Softer shadows can be achieved by blurring and averaging the light-view depth image. We compute the shadowing probability using Chebyshev’s inequality [14] and store the results in a shadow map. Finally, we composite the rendered image with the shadow map to achieve nearly-soft dynamic shadows.
Spray, foam, and bubble
To enhance the realism of fluids, foam, spray and bubble particles are synthesized with [23] as a post-processing step. Fluid-air mixtures are generated at the crest of the wave and in the impact zone of the wave. Spray, foam, and bubble particles are advected by the fluid and dissipate within their predefined lifetimes.
We splat these particles into a foam intensity image using modified additive splatting. Different types of particles use different kernels during splatting [1]. We preferred a larger overall intensity for surface foam particles to increase their visibility, while we used a comparatively smaller intensity value for spray particles to make them less prominent. Furthermore, we used hollow circle structures for the bubble particles to make their appearance more convincing underwater. The kernel typically has a radius of 2 pixels. However, in practice, we found that the kernel radius should be scaled based on the particle’s depth in the view, as particles near the camera occupy more of the view compared to those farther away. Finally, we apply a curve [1] on the foam intensity image to scale it into $[0,1]$ and compose it with rendering.
10 Implementation Details
We set the simulation time step as $0.005$ seconds throughout the simulations. In our PBD solver, we used $10$ iterations for fluids and $50$ iterations for solids for our experiments, since mass particles on the solid models are more strongly coupled than the ones in the fluid. During the PBF simulation, the surface particles of fluids are updated every two time steps.
11 More Results
We show the results of Chair (Fig. 12), Waves (Fig. 13), Garden (Fig. 14), Lego (Fig. 15), Cup & dog (Fig. 16), Headset (Fig. 17), Can (Fig. 18), Astronaut (Fig. 19), Ficus (Fig. 20), and Bulldozers (Fig. 21). For better visualization, please refer to the supplementary video.
@@ 11 More Results
Figure 12: Chair.  A soft chair falls into the pool, causing deformation and ripples.
Figure 13: Waves crashing on cliff. A coastal cliff rises from the sandy beach, while the sea waves continuously crash against the rocky surface, generating splashes and foam upon collision.
@@
Figure 14: Flooding garden.  Water leaks into the garden and submerge the table. As the water level goes up, the surface gets more vibrant and washes the potted plant away.
Figure 15: Splashing LEGO.  Through the two-way coupling dynamics, the LEGO bulldozer is animated to surf on the splashing waves.
@@
Figure 16: “Everything is water”.  Pouring water into the paper cup on the table and transforming the cup and the dog toy into water. The water spills out.
Figure 17: Headset waterfall. Water flows from a headset hanging above an office desk, resembling a faucet. Due to surface tension, the water forms droplets as it falls, sliding down the computer screen and splashing onto the desk, creating a puddle.
Figure 18: Water droplets on can.  Droplets of water fall onto the surface of a soda can, coalesce due to surface tension and gradually overflow.
@@
Figure 19: Black magic.  An astronaut in space strucked by the black magic of the Trisolarans, and get transformed into a water sphere.
Figure 20: Deformable ficus. A deformable ficus plant undergoes continuous shape changes as it is dragged and manipulated by external forces.
Figure 21: LEGO bulldozers in glass bowl. A collection of LEGO bulldozer rigid bodies fall into a round glass box, colliding with each other. They cast shadow on the ground and eventually stack and scatter throughout the box.